OpenStudy (anonymous):

Give solutions of the nonlinear system of equations: y=x^2-12x+36 and x +y=8

OpenStudy (anonymous):

x+y = 8 y = x-8 y = x^2-12x+36 = x-8 x^2-13x+44 = 0 Then use the quadratic formula :D